What multiplies to get -0.25 and adds to get -1
Answer by jorel1380(3719)

Let the two numbers be m and n. Then:
mn=-0.25 and m+n=-1
So:
m=-1-n
(-1-n)n=-0.25
-n-nē=-0.25
nē+n-0.25=0
Using the quadratic formula, we get two roots: 0.207106781187 and -1.20710678119, as m and n. ☺☺☺☺